OpenStudy (anonymous):

what is 2/5 of 20 kg

OpenStudy (anonymous):

20/5=4 4*2=8 8kg

OpenStudy (texaschic101):

2/5 of 20 just means 2/5 * 20/1

OpenStudy (anonymous):

(2 / 5) * (20 kilograms) = 8 kilograms